please help me solve this problem
category:coins and money
1.Alice is counting a cash register at the end of the night.she has $1015 in $5 bills and $10 bills.if she has 138 bills all together,how many of each type does she has?
please help me right away.

Answer by checkley77(12844)

f+t=138 or f=138-t
5(138-t)+10t=1015
690-5t+10t=1015
5t=1015-690
5t=325
t=325/5
t=65 ten dollar bills.
f=138-65=73 ten dollar bills.
Proof:
5*73+10*65=1015
365+650=1015
1015=1015